# Quantum ergodicity and scrambling in quantum annealers

### v1

@@ IV.1 Forward ramps
The evolution along the forward ramp is defined iterativelly via a product of unitaries in time succession, see Eq. (9). For $t=m\Delta t\in[0,T]$, we write $U(t=m\Delta t)=\left[\prod_{p=0}^{m-1}e^{-ip\Delta tH(s(p\delta t))}\right]e^{-im\Delta tH(s(m\Delta t))}$ where the product inside the square brackets is the unitary up to the preceding time $U((m-1)\Delta t)$. The unitary of the full ramp is obtain when the number of steps equals $T/\Delta t$. The spectral decomposition of any of the unitaries along the ramp reads $U(t)|\mu_{l}(t)\rangle=e^{-i\mu_{l}(t)}|\mu_{l}(t)\rangle$, with $\{\mu_{l}(t)\}_{l=1,..,2^{N}}$ the set of eigenphases and $\{|\mu_{l}(t)\rangle\}_{l=1,..,2^{N}}$ the set of associated eigenvectors, and with $t\rightarrow T$ for the unitary of the full evolution $U(T)$.
@@ IV.2 Cyclic ramps
The onset of quantum chaos becomes more subtle (and thus, more interesting) when we consider the evolution along the cyclic ramp. Since the cyclic ramp is composed of a forward and a backward ramps, the first half of the evolution is always described by $U(T)$. Hence, the results of the MLSR study discussed in Sec. IV.1 apply here for the first half of the evolution. For the second half of the evolution completing the cyclic ramp, the results of the MLSR analysis are shown in the same Fig. 4a, for the regime $T\leq t\leq 2T$. We notice that for most of the backward ramp the chaotic properties remain and only for the final portion, $s(t)\gtrapprox 1.9$, we observe a deviation from the COE value, with the MLSR of the eigenphases of $U(2T)$ acquiring an intermediate value between $\overline{r}_{\rm Int}$ and $\overline{r}_{\rm COE}$. Thus, while there is still some degree of level repulsion, the system no longer resembles a fully random unitary. In fact, subsets of the eigenphases can unambigously be labeled as low- and high-energy, and bulk. This means that the spectral structure of $U(2T)$ is more similar to that of a Hamiltonian system, than of a random unitary. We devote the rest of this section to a judicious quantitative study supporting our previous claim.
Recall that for a time-independent many-body Hamiltonian the entanglement entropy $S_{A}$ of eigenstates allows us to distinguish ergodic ($S_{A}$ saturating the Page value) from nonergodic ($S_{A}$ constant) eigenstates. Typically a quantum chaotic Hamiltonian has eigenstates of both types, with nonergodic ones constituting a small fraction appearing close to the lower and top ends of the energy spectrum, and ergodic ones constituting the dominant fraction which cluster close to the middle of the energy spectrum [13].
@@
We thus focus on $U(2T)$ and ask: how similar is the structure of its egenstates $\{|\mu_{l}(2T)\rangle\}_{l}$ as seen in the plane $(\langle H_{\rm M}\rangle,S_{A})$ to that of a chaotic Hamiltonian? Here $H_{\rm M}=H(0)=H(2T)$ defines the natural energy scale. In Fig. 4c we show $S_{A}$ of all the $\{|\mu_{l}(2T)\rangle\}_{l}$ as function of their mean energy $\langle H_{\rm M}\rangle$ for a system of $N=14$ spins. This way of displaying $S_{A}$ unveils a structure with strike similarity to that of a nonintegrable (quantum chaotic) Hamiltonian [13], in stark contrast with the expectation that the cyclic ramp must recover the structure of the integrable initial Hamiltonian $H_{\rm M}$. Further, the structure seen in Fig. 4c shows clear emergent energy sectors which can be classified into low-energy, bulk, and high-energy, distinctions which are absent in the forward-ramp unitary.
@@
The lack of adiabaticity in the bulk of the spectrum is evident in the preceding analysis. A system evolving in perfect adiabatic fashion should connect every eigenstate of $H_{\rm M}$ to itself after the cyclic ramp. Thus, in the $S_{A}$ vs energy diagram, one would observe that only discrete values of $\langle H_{\rm M}\rangle$ are present (in this case, this would be $N+1$ distinct eigenvalues). In contrast to this, we find the values of $\langle H_{\rm M}\rangle$ appear to be continuous in value, particularly near $\langle H_{\rm M}\rangle\approx 0$.
@@
To complete our investigation of the structure of eigenstates of $U(2T)$, we look into the behavior of $S_{A}$ as a function of the system size $N$. In Fig. 5a we show $S_{A}$ vs $\langle H_{\rm M}\rangle$ for all the eigenstates of $U(2T)$ for systems with $N=8,10,12,14$ (purples, greens, blues, black, respectively). We find that, even for the smallest size considered, signatures of a structure resembling that of a nonintegrable Hamiltonian are noticeable. We thus study $S_{A}(N)$ in terms of an energy density $E_{\rm d}=\langle H_{\rm M}\rangle/N\in[-1,1]$. We bin the $E_{\rm d}$ axis into small energy-density windows $E_{\rm W}=[E_{\rm d}-\Delta E_{\rm d},E_{\rm d}+\Delta E_{\rm d}]$, whose size is chosen as to not have empty windows. For each system size we count the number of eigenstates inside a given energy-density window $N_{\rm W}=|\{S_{A}(\epsilon):\epsilon\in E_{\rm W}\}|$ with |.||.|| . | indicating the cardinality of the set, and compute the energy-window averaged $S_{A}$, that is $\left.\overline{S_{A}}\right|_{E_{\rm W}}=\frac{1}{N_{W}}\sum_{\epsilon\in E_{\rm W}}S_{A}(\epsilon)$. Finally we regress the energy-density window averaged $S_{A}$’s to a linear model, that is, $\left.\overline{S_{A}}\right|_{E_{\rm W}}(N)\leftarrow aN+b$ with $(a,b)$ the regression coefficients.
@@
In Fig. 5b we show the values of $a$ as a function of $E_{\rm d}$. We recognize two distinct behaviors, volume-law entanglement and area-law entanglement, with the particularity that the volume-law eigenstates can be subdivided into two different groups. For $|E_{\rm d}|<0.5$ we can unambiguously conclude that the eigenstates of $U(2T)$ obey a volume-law of entanglement which is consistent with the Page result for random pure states, and that the majority of eigenstates are found in this region. For $0.5<|E_{\rm d}|\lessapprox 0.8$, we encounter few sets of eigenstates which do follow a volume-law of entanglement but whose values of $(a,b)$ indicate they might be random Gaussian states [40]. Finally, for $|E_{\rm d}|>0.8$ we observe four sets of eigenstates which obey an area-law of entanglement. Interestingly for all system sizes studied, we can readily recognize the states in this last category. Naturally, the lowest and highest energy states, at $E_{\rm d}=\pm 1$, correspond to the ground state $|-\rangle^{\otimes N}$ and highest excited state $|+\rangle^{\otimes N}$ of $H_{\rm M}$, which appear both in Fig. 4c and Fig. 5a at $S_{A}=0$ (first black dashed line). The next one, in increasing $S_{A}$, corresponds to Dicke states with one flip (low-energy) and $N-1$ flips (high-energy), and states with similar entanglement structure, appearing at $S_{A}=\log(2)$ (orange dashed line) in Fig. 4c and Fig. 5a.
@@ V Operator dynamics and information scrambling
$P_{k}(t)=\frac{1}{d^{2}\>\mathrm{Tr}\left(A^{2}\right)}\sum\limits_{r(Q)=k}\lvert\mathrm{Tr}\left(Q\>U(t)^{\dagger}AU(t)\right)\rvert|^{2},$ (13)
@@
Evidently, $1\leq\mu(t)\leq N$. For local operators, one expects $\mu(t)$ to grow at short times (in some cases, exponentially [47]) and, under generic dynamics, the mean operator size becomes extensive, i.e., proportional to the system size $N$. Particularly, a Haar-random unitary leads to $\mu_{\rm Haar}\simeq 3N/4$ independent of our choice of initial Pauli operator. We point out that the properties of the mean operator size are tightly connected with that of out-of-time-ordered correlators, as discussed in Ref. [46]. Naturally, higher order moments of the OSD can also be considered as more refined descriptions of the scrambling process [43, 46].
@@ V.2 Dynamics of global operators
As we discussed above, the operator dynamics for the set of single-site Pauli operators shows clear signatures of information scrambling. In particular, we found that the mean operator size dynamics is equivalent to that generated by Haar-random dynamics, meaning that the “gentle” adiabatic drive has a similar effect as one would expect from a thermalizing periodically driven system, or even from a random quantum circuit. However, as we have seen in Sec. IV, the cyclic adiabatic evolution is non-generic and deviations from chaotic evolution are expected. Here we discuss such cases for the operator dynamics.
@@
In Fig. 6 (d)-(f) we display the mean operator size for the global operators $\{S_{x},S_{y},S_{z}\}$. The dynamics bears similarities and differences with the evolution of single-site operators. For the forward ramp, $0\leq t\leq T$, we observe that operator sizes become clearly extensive with values matching those of random operators for $S_{x}$ and $S_{y}$; the case of $S_{z}$ shows maximum operator size increasing with system size, but clearly deviates from the Haar prediction. From all cases considered, the one that stands out the most is $S_{x}$, where we observe that size of $S_{x}$ shrinks back to $\sim 1$ during the backward ramp. Thus, we see that the dynamics of $S_{x}$ reverses back to its initial configuration, although not perfectly. A key aspect here is that $S_{x}$ is a special operator in this problem, since its proportional to the mixer Hamiltonian, i.e. $H(0)=H(2T)=H_{M}=2S_{x}$. At first sight, the evolution of $S_{x}$ seen here is consistent with the expectation that the adiabatic evolution should connect the eigenstates of $H_{M}$ back to themselves at the end of the cyclic protocol. However, a more detailed inspection reveals that only the eigenstates of $S_{x}$ near the edges of its spectrum follow this expected evolution, while eigenstates near the bulk undergo mixing evolution. We show this analysis in Appendix A. As evidenced from the results in this section, it suffices that the dominant eigenstates (i.e. those with larger eigenvalues in absolute value) evolve adiabatically for the operator size to shrink back to its original configuration.
@@ VI Discussion and outlook
In this work we have shown that a basic model of a quantum annealer can display clear signatures of quantum ergodicity and scrambling. These features become prominent when one considers the dynamical behavior of these models beyond the low (and high) energy regime. Here, we have analyzed three specific aspects which fall into this criterion: the generation of volume-law entanglement when the system is initialized in a product state which is rotated from the ground state, the random-matrix-like spectral properties of the complete unitary map $U(T)$ describing the annealing dynamics, and the onset of information scrambling as measured by the growth of generic operators in the Heisenberg picture. Furthermore, we have shown that the counterintuitive fact that slow adiabatic drive results in strongly chaotic dynamics is balanced by the existence of clear deviations from chaoticity which are observed when a cyclic adiabatic drive is considered.
@@
Beyond its relevance for quantum annealing devices, our study is naturally connected to fundamental aspects of nonequilibrium many-body quantum systems under adiabatic driving. While here we have studied a model defined on a chaotic path (see Fig. 1a), setting $\lambda=0$ in Eq. (1) would actually render the model integrable for all values of $s$. In addition, adiabaticity in the bulk of the spectrum could be easier to achieve in such case thanks to the existence of conserved quantities leading to exact crossings in the spectrum. Recent studies have focused on a complimentary model featuring all-to-all interactions, but where the addition of driving leads to quantum chaos signatures [32]. In this vein, we notice that a small $\lambda>0$ (i.e., a weak integrability breaking) leads to avoided crossings with very small gaps, in fact smaller than in the chaotic regime, which displays level repulsion. The characterization of integrability-to-chaos transition in terms of adiabatic process has also been studied recently [52, 53], and integrable many-body adiabatic dynamics has been shown to impact the proliferation of errors in quantum annealing [34]. In this context, further characterizing the interplay between adiabaticity and (non) integrability in a more general setting stands out as an interesting open problem.

### v2

@@ IV.1 Forward ramps
The evolution along the forward ramp is defined iterativelly via a product of unitaries in time succession, see Eq. (9). For $t=m\Delta t\in[0,T]$, we write $U(t=m\Delta t)=\left[\prod_{p=0}^{m-1}e^{-ip\Delta tH(s(p\Delta t))}\right]e^{-im\Delta tH(s(m\Delta t))}$ where the product inside the square brackets is the unitary up to the preceding time $U((m-1)\Delta t)$. The unitary of the full ramp is obtained when the number of steps equals $T/\Delta t$. The spectral decomposition of any of the unitaries along the ramp reads $U(t)|\mu_{l}(t)\rangle=e^{-i\mu_{l}(t)}|\mu_{l}(t)\rangle$, with $\{\mu_{l}(t)\}_{l=1,..,2^{N}}$ the set of eigenphases and $\{|\mu_{l}(t)\rangle\}_{l=1,..,2^{N}}$ the set of associated eigenvectors, and with $t\rightarrow T$ for the unitary of the full evolution $U(T)$.
@@ IV.2 Cyclic ramps
The onset of quantum chaos becomes more subtle (and thus, more interesting) when we consider the evolution along the cyclic ramp. Since the cyclic ramp is composed of a forward and a backward ramps, the first half of the evolution is always described by $U(T)$. Hence, the results of the MLSR study discussed in Sec. IV.1 apply here for the first half of the evolution. For the second half of the evolution completing the cyclic ramp, the results of the MLSR analysis are shown in the same Fig. 4a, for the regime $T\leq t\leq 2T$. We notice that for most of the backward ramp the chaotic properties remain and only for the final portion, $s(t)\gtrapprox 1.9$, we observe a deviation from the COE value, with the MLSR of the eigenphases of $U(2T)$ acquiring an intermediate value between $\overline{r}_{\rm Int}$ and $\overline{r}_{\rm COE}$. Thus, while there is still some degree of level repulsion, the system no longer resembles a fully random unitary. In fact, there exists a choice of Hamiltonian with respect to which the mean energies of eigenstates together with their entanglement entropies $S_{A}$ allows us to consistently divide the spectrum into sectors of low- and high-energy, and bulk. For us the choice of reference Hamiltonian is informed by the structure of the cyclic ramp. This means that the spectral structure of $U(2T)$ is more similar to that a chaotic Hamiltonian system, than of a random unitary. We devote the rest of this section to a judicious quantitative study supporting our previous claim.
The entanglement entropy $S_{A}$ of eigenstates allows us to distinguish ergodic ($S_{A}$ saturating the Page value) from nonergodic ($S_{A}$ constant) eigenstates. Generic many-body Hamiltonians have eigenstates of both classes. For integrable and near-integrable Hamiltonians, ergodic eigenstates, if any, are found at energies close to the middle of the spectrum, while nonergodic eigenstates may be found at all energies [44]. For a nonintegrable chaotic Hamiltonian the two types of eigenstates occupy distinct energy ranges. In fact, as $E\to\overline{E}$ with $\overline{E}={\rm tr}(H)/2^{N}$ the mean energy, from above or below, a crossover between $S_{A}\propto N^{0}$ and $S_{A}\propto N^{1}$ takes place [45], indicating that nonergodic (area-law) eigenstates only exist at the edges of the spectrum (with some notable exceptions [46]). As such, $S_{A}$ as a function of $E$ for the eigenstate of many-body Hamiltonians is a concave function of the energies with a maximum at the “middle”, i.e., the mean energy $\overline{E}$, with a characteristic “arch” shape. Importantly, for chaotic Hamiltonians the average and variance of $S_{A}$ on only the mid-spectrum eigenstates follows well characterized behaviors [47, 48, 13].
@@
We thus focus on $U(2T)$ and ask: how similar is the structure of its egenstates $\{|\mu_{l}(2T)\rangle\}_{l}$ as seen in the plane $(\langle H_{\rm M}\rangle,S_{A})$ to that of a chaotic Hamiltonian? Here $H_{\rm M}=H(0)=H(2T)$ defines the natural energy scale. In Fig. 4c we show $S_{A}$ vs $\langle H_{\rm M}\rangle$ for all the $\{|\mu_{l}(2T)\rangle\}$ for a system of $N=14$ spins. Clearly the data follows a curve with all the characteristics of the $S_{A}$ vs $E$ curve of a chaotic Hamiltonian described before. In particular, we observe a concave shape with a maximum at the mean energy; we provide a semi analytical form for this function in App. B, together with an analysis of the expected behavior of the mean and variance of $S_{A}$ for the “mid-spectrum” eigenstates (see App. B for details). Further, the structure seen in Fig. 4c shows clear emergent energy sectors which can be classified into low-energy, bulk, and high-energy, distinctions which are absent in the forward-ramp unitary.
Physically, then, we observe sharp difference between evolutions stemming from the the forward and cyclic annealing schedule. While the forward dynamics is described by a unitary process which is fully chaotic across all energy scales, the cyclic dynamics shows an emergent structure in which low- and high-energy eigenstates states deviate substantially from the typical chaotic predictions. As we will confirm in the next section, this deviation can be linked to adiabaticity in those parts of the energy spectrum.
@@
We also note that the lack of adiabaticity in the bulk of the spectrum is evident in the preceding analysis. A system evolving in perfect adiabatic fashion should connect every eigenstate of $H_{\rm M}$ to itself after the cyclic ramp. Thus, in the $S_{A}$ vs energy diagram, one would observe that only discrete values of $\langle H_{\rm M}\rangle$ are present (in this case, this would be $N+1$ distinct eigenvalues). In contrast to this, we find the values of $\langle H_{\rm M}\rangle$ appear to be continuous in value, particularly near $\langle H_{\rm M}\rangle\approx 0$.
@@
To complete our investigation of the structure of eigenstates of $U(2T)$, we look into the behavior of $S_{A}$ as a function of the system size $N$. In Fig. 5a we show $S_{A}$ vs $\langle H_{\rm M}\rangle$ for all the eigenstates of $U(2T)$ for systems with $N=8,10,12,14$ (purples, greens, blues, black, respectively). We find that, even for the smallest size considered, signatures of a structure resembling that of a nonintegrable Hamiltonian are noticeable. We thus study $S_{A}(N)$ in terms of an energy density $E_{\rm d}=\langle H_{\rm M}\rangle/N\in[-1,1]$. We bin the $E_{\rm d}$ axis into small energy-density windows $E_{\rm W}=[E_{\rm d}-\Delta E_{\rm d},E_{\rm d}+\Delta E_{\rm d}]$, whose size is chosen as to not have empty windows. For each system size we count the number of eigenstates inside a given energy-density window $N_{\rm W}=|\{S_{A}(\epsilon):\epsilon\in E_{\rm W}\}|$ with $|.|$ indicating the cardinality of the set, and compute the energy-window averaged $S_{A}$, that is $\left.\overline{S_{A}}\right|_{E_{\rm W}}=\frac{1}{N_{W}}\sum_{\epsilon\in E_{\rm W}}S_{A}(\epsilon)$. Finally we regress the energy-density window averaged $S_{A}$’s to a linear model, that is, $\left.\overline{S_{A}}\right|_{E_{\rm W}}(N)\leftarrow aN+b$ with $(a,b)$ the regression coefficients.
@@
In Fig. 5b we show the values of $a$ as a function of $E_{\rm d}$. We recognize two distinct behaviors, volume-law entanglement and area-law entanglement, with the particularity that the volume-law eigenstates can be subdivided into two different groups. For $|E_{\rm d}|<0.5$ we can unambiguously conclude that the eigenstates of $U(2T)$ obey a volume-law of entanglement which is consistent with the Page result for random pure states, and that the majority of eigenstates are found in this region. For $0.5<|E_{\rm d}|\lessapprox 0.8$, we encounter few sets of eigenstates which do follow a volume-law of entanglement but whose values of $(a,b)$ indicate they might be random Gaussian states [41]. Finally, for $|E_{\rm d}|>0.8$ we observe four sets of eigenstates which obey an area-law of entanglement. Interestingly for all system sizes studied, we can readily recognize the states in this last category. Naturally, the lowest and highest energy states, at $E_{\rm d}=\pm 1$, correspond to the ground state $|-\rangle^{\otimes N}$ and highest excited state $|+\rangle^{\otimes N}$ of $H_{\rm M}$, which appear both in Fig. 4c and Fig. 5a at $S_{A}=0$ (first black dashed line). The next one, in increasing $S_{A}$, corresponds to Dicke states with one flip (low-energy) and N1N-1 flips (high-energy), and states with similar entanglement structure, appearing at $S_{A}=\log(2)$ (orange dashed line) in Fig. 4c and Fig. 5a.
@@ V Operator dynamics and information scrambling
$P_{k}(t)=\frac{1}{d^{2}\>\mathrm{Tr}\left(A^{2}\right)}\sum\limits_{r(Q)=k}\lvert\mathrm{Tr}\left(Q\>U(t)^{\dagger}AU(t)\right)\rvert^{2},$ (13)
@@
Evidently, $1\leq\mu(t)\leq N$. For local operators, one expects $\mu(t)$ to grow at short times (in some cases, exponentially [53]) and, under generic dynamics, the mean operator size becomes extensive, i.e., proportional to the system size $N$. Particularly, a Haar-random unitary leads to $\mu_{\rm Haar}\simeq 3N/4$ independent of our choice of initial Pauli operator. Naturally, higher order moments of the OSD can also be considered as more refined descriptions of the scrambling process [49, 52].
@@ V.2 Dynamics of global operators
As we discussed above, the operator dynamics for the set of single-site Pauli operators shows clear signatures of information scrambling. In particular, we found that the mean operator size dynamics is equivalent to that generated by Haar-random dynamics, meaning that the “gentle” adiabatic drive has a similar effect as one would expect from a thermalizing periodically driven system, or even from a random quantum circuit. However, as we have seen in Sec. IV, the evolution stemming from the cyclic path showed clear deviations from chaotic behavior. Here we investigate such behavior for the case of operator dynamics.
In Fig. 6 (d)-(f) we display the mean operator size for the global operators $\{S_{x},S_{y},S_{z}\}$. The dynamics bears similarities and differences with the evolution of single-site operators. For the forward ramp, $0\leq t\leq T$, we observe that operator sizes become clearly extensive with values matching those of random operators for $S_{x}$ and $S_{y}$; the case of $S_{z}$ shows maximum operator size increasing with system size, but clearly deviates from the Haar prediction. We attribute this to the fact that the operator commutes with the interacting Hamiltonian $H_{P}$ (notice that a similar, but weaker effect, is observed for the local Pauli $\sigma_{z}$), which is responsible for the scrambling behavior.
@@
From all cases considered, the one that stands out the most is $S_{x}$, where we observe that size of $S_{x}$ shrinks back to $\sim 1$ during the backward ramp. Thus, we see that the dynamics of $S_{x}$ reverses back to its initial configuration, although not perfectly. A key aspect here is that $S_{x}$ is a special operator in this problem, since its proportional to the mixer Hamiltonian, i.e. $H(0)=H(2T)=H_{M}=2S_{x}$. At first sight, the evolution of $S_{x}$ seen here is consistent with the expectation that the adiabatic evolution should connect the eigenstates of $H_{M}$ back to themselves at the end of the cyclic protocol. However, a more detailed inspection reveals that only the eigenstates of $S_{x}$ near the edges of its spectrum follow this expected evolution, while eigenstates near the bulk undergo mixing evolution, due to imperfect adiabaticity in that regime. We show this analysis in Appendix A. As evidenced from the results in this section, it suffices that the dominant eigenstates (i.e. those with larger eigenvalues in absolute value) evolve adiabatically for the operator size to shrink back to its original configuration. Thus, in line with the results obtained in Sec. IV.2, we find a non-trivial balance between strong, random-like chaotic behavior (e.g. as seen in the the behavior of $S_{y}$), and clear signatures of the near-adiabatic evolution leading to non-generic dynamics (e.g. the operator shrinking of $S_{x}$), including cases exhibiting a mixture of both phenomena (like $S_{z}$).
V.3 Connection to out-of-time ordered correlators and route to experimental exploration
Witnessing chaotic signatures in quantum many-body systems is a challenging task for experiments. This is due to the fact that many measures of chaoticity are computed over many-body eigenstates, which are typically hard to prepare. Even for those states which can be accessed (for instance following a quench), typical local correlation functions struggle to witness chaotic behavior, and more intricate measures are needed. This is evidenced from our own analysis up to here, where we needed to compute quantities like entanglement entropies, unitary eigenphases, or operator wavefunctions.
Nonetheless, significant attention has been given to so-called out-of-time-ordered correlators (OTOCs) as a way to probe chaotic behavior in experiments. These quantities take the form
$C_{A,B}(t)=\mathrm{Tr}\left(A(t)^{\dagger}B^{\dagger}A(t)B\right)$ (17)
where $A$ and $B$ are general operators and $A(t)=U(t)^{\dagger}AU(t)$. Many proposals have been put forward in the past decade to measure OTOCs leveraging different experimental capabilities [54, 55, 56, 57], and recently many of them have been successfully demonstrated in, e.g., superconducting qubits [58], Rydberg atoms [59], and trapped ions [60]. Importantly for our discussion, OTOCs have a direct connection to the mean operator size $\mu(t)$; as this can be recovered by considering the operator $B$ in Eq. (17) to be single-site Paulis. More specifically, one has that [49, 52]
$\mu(t)=\frac{3N}{4}\left(1-\mathcal{M}(t)\right),$ (18)
where
$\mathcal{M}(t)=\frac{1}{3N}\sum\limits_{r(B)=1}C_{A,B}(t)$ (19)
i.e. an average over all OTOCs with single-site Paulis. Thus, we see that the scrambling behavior analyzed in this section has a route for potential experimental exploration, which is rooted in the extensive studies of OTOCs in various platforms.
@@ VI Discussion and outlook
In this work we have shown that a basic model of a quantum annealer can display clear signatures of quantum ergodicity and scrambling. These features become prominent when one considers the dynamical behavior of these models beyond the low (and high) energy regime. Here, we have analyzed three specific aspects which fall into this criterion: the generation of volume-law entanglement when the system is initialized in a product state which is rotated from the ground state, the random-matrix-like spectral properties of the complete unitary map $U(T)$ describing the annealing dynamics, and the onset of information scrambling as measured by the growth of generic operators in the Heisenberg picture. Furthermore, we have shown that the counterintuitive fact that slow adiabatic drive results in strongly chaotic dynamics is balanced by the existence of clear deviations from chaoticity which are observed when a cyclic adiabatic drive is considered. In this regime, we found that adiabaticity plays a major role in explaining such deviations. A key example of this is the observation of reverse scrambling in the Heisenberg dynamics of certain collective operators.
@@
Beyond its relevance for quantum annealing devices, our study is naturally connected to fundamental aspects of nonequilibrium many-body quantum systems under adiabatic driving. A straightforward extension of our work would be to study the dependence of the chaos signatures on the rate of adiabatic driving, i.e. $\dot{s}(t)$. Particularly, a faster sweep of $s(t)$ would bring the dynamics closer to a standard ‘quench’ where the system is changed quickly between two noncommuting (static) Hamiltonians. The deviation from adiabaticity is expected to lead to stronger mixing; however, we have shown here that the adiabatic regime shows features typically connected to driven systems (like the CUE-like spectra). Thus, we believe a nontrivial balance between adiabaticity and chaos would be accessible in such regime.
Another interesting path for future exploration is the assessment of how integrability of $H(s)$ would affect the results shown here. While here we have studied a model defined on a chaotic path (see Fig. 1a), setting $\lambda=0$ in Eq. (1) would actually render the model integrable for all values of $s$. In such case, preliminary results indicate that adiabaticity in the bulk of the spectrum could be easier to achieve thanks to the existence of conserved quantities leading to exact crossings in the spectrum. Recent studies have focused on a complimentary model featuring all-to-all interactions, but where the addition of driving leads to quantum chaos signatures [32]. In this vein, we notice that a small $\lambda>0$ (i.e., a weak integrability breaking) leads to avoided crossings with very small gaps, in fact smaller than in the chaotic regime, which displays level repulsion. The characterization of integrability-to-chaos transition in terms of adiabatic process has also been studied recently [64, 65], and integrable many-body adiabatic dynamics has been shown to impact the proliferation of errors in quantum annealing [34]. In this context, further characterizing the interplay between adiabaticity and (non) integrability in a more general setting stands out as an interesting open problem.